# Identifying and Analyzing Performance-Critical Tokens in Large Language Models

### v3

@@ Appendix I Analysis of the Next-line Tokens
Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
OpenLlama 3B | Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | 58.0 |
[…]
temp | 41.9 | 87.2 | 26.0 | 56.3 | 58.5 | 57.4 | 5.4 | |
Llama 7B | Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | 74.8 |
stop | 80.4 | 94.6 | 61.1 | 68.0 | 67.2 | 72.0 | 0.9 | |
temp | 64.5 | 84.1 | 54.0 | 58.0 | 56.8 | 54.3 | 12.8 | |
Llama 13B | Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | 70.4 |
stop | 81.2 | 94.1 | 59.3 | 76.9 | 69.2 | 40.6 | 0.2 | |
temp | 74.1 | 80.0 | 46.5 | 30.6 | 58.3 | 25.4 | 17.9 | |
Llama 33B | Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | 81.6 |
stop | 84.3 | 95.6 | 65.7 | 77.6 | 78.6 | 81.8 | 1.0 | |
[…]
Llama 2 7B | Standard ICL | 70.7 | 85.0 | 93.2 | 58.3 | 66.7 | 66.3 | 55.0 |
stop | 85.5 | 92.7 | 56.4 | 66.6 | 63.6 | 57.1 | 0.4 | |
temp | 69.8 | 82.8 | 56.3 | 58.8 | 67.5 | 42.9 | 7.7 | |
Llama 2 13B | Standard ICL | 73.6 | 82.8 | 94.9 | 62.8 | 74.6 | 71.2 | 55.4 |
stop | 81.2 | 94.5 | 61.2 | 73.7 | 72.0 | 53.2 | 1.0 | |
temp | 71.1 | 95.6 | 61.0 | 72.4 | 72.9 | 54.3 | 2.4 | |
Mistral 7B | Standard ICL | 80.2 | 82.2 | 97.0 | 67.4 | 82.4 | 73.6 | 78.6 |
stop | 81.2 | 97.3 | 65.5 | 82.0 | 77.6 | 73.9 | 0.6 | |
@@
Appendix J Comparison Experiments with More Text Examples
In the ideal scenario, our experiments would have been conducted on the full test set. However, in practice, this is infeasible for any of the models studied in our paper due to computational resource constraints. For instance, it took 42 hours for the OpenLlama 3B model to run one round of the representation ablation experiment on the whole test set of DBPedia (i.e., one cell in Table 2 and 3 for the DBPedia column). To verify our number of test examples decision, we provide additional results where we scale up the number of test examples and observe no difference with our original experimental setup. Thus, we believe that limiting our test set sample size to 500 is a reasonable setup. We provide the test set statistics and the experiment results in Table 13 and Table 14.
For TREC, RTE, and CB, using 500 test examples won’t affect the final results at all since their test set size is smaller than 500. We provide the results of experiments using all test examples in SST2, and 5000 test examples in AGNews and DBPedia here to prove our point that limiting our test set sample size to 500 is a reasonable compromise. Shown in Table 14, compared to the results we show in Table 2 and Table 3, the numbers are changed less than 1% for all the results.
Appendix K Results using different instruction prompts
@@
Appendix L Representation-level ablation on Machine Translation Tasks
Besides the classification tasks, we also show results in the machine translation tasks to show that our findings could also be extended in text generation tasks. We used the Flores MT dataset (Costa-jussà et al., 2022) to conduct this set of 4-shot machine translation experiments. The results are reported with the BLEU metric (Papineni et al., 2002). We investigated three different language directions: English-to-French, English-to-Danish, and English-to-German. We used 10 random seeds for En-Fr and En-De and 15 random seeds for En-Da to randomly choose the demonstrations. 100 test examples are sampled in this set of the experiments as a computational compromise. Similar to the classification tasks, we keep the answer (i.e., target language) unablated for all the settings and ablate different kinds of tokens. Results in Table 23 show the consistent finding to those in Section 5.2.1.
Appendix M Representation-level ablation on Question Answering Tasks
We show the representation-level experimental results of the question answering (QA) tasks in this section. We used Commonsense QA (Talmor et al., 2019) dataset to test if the template and stopword tokens would directly affect the downstream task performance. We applied the settings of 4 in-context examples and 15 random seeds in this set of experiments. We frame the task as directly answering the questions instead of choosing one answer from the choices because the token types in this scenario are easier to be categorized.
@@
Setting | OpenLlama 3B | Llama 7B | Llama 13B | Llama 33B | Llama 2 7B | Llama 2 13B | Mistral 7B |
Zero-shotcont | 7.42 | 16.62 | 14.49 | 19.47 | 17.51 | 17.78 | 16.64 |
Zero-shotstop | 11.71 | 21.96 | 18.98 | 23.38 | 22.13 | 22.31 | 19.16 |
Zero-shottemp | 13.24 | 24.38 | 25.73 | 27.20 | 25.42 | 25.11 | 25.56 |
Standard ICLcont | 14.40 | 24.07 | 26.22 | 27.73 | 25.89 | 24.71 | 25.47 |
Standard ICLstop | 11.96 | 21.84 | 21.44 | 26.93 | 24.62 | 23.09 | 23.69 |
@@
Appendix N Representation-level Ablation based on the Token Count
@@
Tokenizer | Setting | Avg. | AGNews | DBPedia | SST2 | TREC | CB | RTE |
OpenLlama 3B | cont | 204.1 | 207.5 | 278.8 | 116.3 | 48.7 | 295.5 | 278.0 |
stop | 43.0 | 43.2 | 45.1 | 27.7 | 21.8 | 66.7 | 53.7 | |
temp | 56.5 | 43.3 | 49.9 | 42.4 | 48.8 | 78.5 | 76.3 | |
Llama & Llama 2 | cont | 220.6 | 238.5 | 288.8 | 127.8 | 51.3 | 312.0 | 305.1 |
stop | 45.0 | 43.1 | 46.1 | 29.9 | 21.7 | 71.1 | 57.9 | |
temp | 52.7 | 41.3 | 50.5 | 48.7 | 36.7 | 70.9 | 68.1 | |
Mistral | cont | 213.6 | 225.0 | 284.7 | 123.7 | 50.1 | 304.2 | 293.7 |
stop | 44.7 | 43.2 | 45.0 | 29.9 | 21.7 | 70.7 | 57.8 | |
@@
Models | Setting | AGNews | DBPedia | SST2 | TREC | CB | RTE | Avg. |
10 Random Tokens | ||||||||
Llama 2 7B | From all | 14.5 | 16.9 | 71.7 | 47.3 | 60.2 | 57.5 | 44.7 |
From cont | 3.0 | 9.2 | 60.9 | 61.9 | 65.2 | 60.4 | 43.4 | |
From stop | 15.2 | 35.6 | 75.2 | 68.5 | 62.0 | 64.1 | 53.4 | |
From temp | 53.0 | 50.1 | 59.1 | 56.2 | 64.4 | 60.2 | 57.2 | |
Mistral 7B | From all | 67.8 | 67.2 | 75.0 | 70.9 | 63.8 | 60.7 | 67.6 |
From cont | 36.6 | 64.0 | 63.1 | 67.5 | 59.3 | 57.3 | 58.0 | |
From stop | 73.5 | 68.6 | 86.3 | 72.8 | 61.8 | 61.6 | 70.8 | |
From temp | 78.7 | 78.7 | 92.1 | 68.9 | 71.2 | 68.4 | 76.3 | |
20 Random Tokens | ||||||||
Llama 2 7B | From all | 12.6 | 24.0 | 77.2 | 54.4 | 60.6 | 58.9 | 47.9 |
From cont | 1.3 | 8.6 | 65.3 | 63.5 | 61.7 | 62.2 | 43.8 | |
From stop | 28.3 | 45.5 | 84.5 | 66.8 | 58.0 | 67.1 | 58.4 | |
From temp | 75.0 | 63.7 | 60.3 | 59.0 | 65.4 | 60.7 | 64.0 | |
Mistral 7B | From all | 78.4 | 67.8 | 74.0 | 70.1 | 65.4 | 62.0 | 69.6 |
From cont | 69.1 | 63.3 | 59.3 | 68.2 | 60.5 | 56.8 | 62.9 | |
From stop | 78.4 | 74.3 | 89.9 | 74.2 | 68.3 | 69.1 | 75.7 | |
From temp | 81.9 | 78.7 | 93.0 | 68.9 | 72.4 | 71.4 | 77.7 | |
30 Random Tokens | ||||||||
Llama 2 7B | From all | 27.4 | 27.7 | 77.9 | 54.9 | 63.9 | 62.2 | 52.3 |
From cont | 1.3 | 7.6 | 77.9 | 66.3 | 64.2 | 63.5 | 46.8 | |
From stop | 44.8 | 59.8 | 92.8 | 68.3 | 53.2 | 69.3 | 64.7 | |
From temp | 76.9 | 66.1 | 61.0 | 59.6 | 67.9 | 59.8 | 65.2 | |
Mistral 7B | From all | 79.9 | 74.6 | 83.0 | 67.8 | 69.0 | 64.6 | 73.2 |
From cont | 71.0 | 62.6 | 58.2 | 67.2 | 59.2 | 59.4 | 62.9 | |
From stop | 79.8 | 76.9 | 91.0 | 74.2 | 71.1 | 68.8 | 77.0 | |
From temp | 84.0 | 79.3 | 93.8 | 68.9 | 75.6 | 75.4 | 79.5 | |
40 Random Tokens | ||||||||
Llama 2 7B | From all | 45.2 | 27.2 | 77.9 | 54.9 | 58.9 | 62.3 | 54.4 |
From cont | 0.9 | 9.1 | 89.5 | 66.1 | 62.3 | 62.4 | 48.4 | |
From stop | 49.2 | 63.7 | 94.0 | 68.7 | 53.7 | 70.8 | 66.7 | |
From temp | 77.9 | 66.1 | 62.9 | 61.5 | 68.1 | 60.1 | 66.1 | |
Mistral 7B | From all | 78.8 | 73.4 | 86.9 | 69.2 | 69.3 | 67.4 | 74.2 |
From cont | 74.1 | 59.4 | 55.8 | 68.0 | 61.0 | 59.9 | 63.0 | |
From stop | 80.1 | 76.4 | 91.0 | 74.2 | 72.0 | 69.6 | 77.2 | |
From temp | 85.0 | 80.7 | 93.8 | 69.0 | 76.4 | 76.2 | 80.2 |
Appendix O Results of the Token-level Ablation
@@
Detailed results of the token-level ablation are shown in Table 27. We omited the temp case from here since it constantly yields an accuracy of 0% when both the template token in the demonstrations and the test examples are ablated. Since the setting for the template tokens are not aligned with the ones for the stopword and content tokens, we included another set of experiments where only the the template tokens in the demonstrations are ablated at the token level in Appendix P. We want to emphasize that this experimental design choice does not affect the main findings in Section 5.2.2, where information is being propagated from the representations of content tokens to the representations of the performance-critical tokens and this incorporation of the information makes them better at affecting task performance, partially explaining the working mechanism of in-context learning.
Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. |
OpenLlama 3B | Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | 58.0 |
cont | 31.5 | 63.0 | 40.6 | 25.4 | 56.1 | 48.9 | 44.3 | |
stop | 64.4 | 91.5 | 20.9 | 62.3 | 57.8 | 52.6 | 58.3 | |
Llama 7B | Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | 74.8 |
cont | 55.2 | 67.2 | 42.6 | 50.8 | 57.4 | 56.3 | 54.9 | |
stop | 82.3 | 93.8 | 64.1 | 69.7 | 66.5 | 70.0 | 74.4 | |
Llama 13B | Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | 70.4 |
cont | 78.8 | 81.7 | 45.3 | 75.1 | 55.1 | 54.5 | 65.1 | |
stop | 82.5 | 92.5 | 61.5 | 76.5 | 69.6 | 40.5 | 70.5 | |
Llama 33B | Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | 81.6 |
cont | 74.0 | 89.6 | 67.0 | 73.0 | 69.8 | 49.0 | 70.4 | |
@@
Appendix P Token-level Ablation for Template Tokens
To maintain consistency of the templates across both demonstrations and test examples, we choose to ablate the template tokens at the token level in both in Section 5.2.2. This experimental design differs from the other two token-level ablations. This inconsistency does not impact the main findings in Section 5.2.2, which show that information is propagated from the representations of content tokens to the representations of performance-critical tokens and this information aggregation enhances the ability of performance-critical tokens to improve the final task performance, partially explaining the mechanism of in-context learning. For completeness, we provide a supplemental experiment in this section where only the template tokens in the demonstrations are ablated.
@@
Models | Settings | Avg. | AGNews | SST2 | TREC | DBPedia | RTE | CB |
OpenLlama 3B | Standard ICL | 58.0 | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 |
temp | 17.4 | 0.0 | 24.2 | 40.5 | 0.4 | 35.6 | 3.6 | |
Llama 7B | Standard ICL | 74.8 | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 |
temp | 29.0 | 41.0 | 11.4 | 39.9 | 0.9 | 49.7 | 31.3 | |
Llama 13B | Standard ICL | 70.4 | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 |
temp | 36.9 | 82.1 | 17.5 | 51.4 | 8.6 | 39.5 | 22.0 | |
Llama 33B | Standard ICL | 81.6 | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 |
@@
Appendix Q Possible Applications
@@
This area is indeed attracting increased research attention, and our findings could contribute valuable insights into ongoing work on efficient sequence modeling and memory management in LLMs (Liu et al., 2023b; Zhang et al., 2023; Bai et al., 2024). By identifying which tokens retain critical task-related information, our work aligns with and can inform methods focused on compressing intermediate states and improving long-context processing for applications ranging from summarization and document understanding to interactive dialogue systems.
@@
Given our findings, there is a promising avenue for improving in-context learning (ICL) performance by including a greater number of examples in ICL prompts (Li et al., 2023a; Hao et al., 2022; Bertsch et al., 2024). Our results suggest that only a subset of token representations, specifically performance-critical tokens, play a critical role in determining ICL performance, while the representations of other tokens are less impactful. This observation opens up the possibility of selectively compressing or omitting unimportant token representations after the initial encoding of a demonstration. By doing so, it becomes feasible to maximize the use of the model’s fixed-length capacity, potentially enabling the inclusion of a higher number of examples within the same prompt length constraints. This approach may enhance the effectiveness of ICL in tasks where the availability of diverse examples contributes to improved model accuracy and stability.
@@
Appendix R Results of Representation-level Partial performance-critical Token Ablation

### v4

@@ Appendix I Stopword Tokens
| Models | Setting | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. | ||
Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | 58.0 | ||
| stop | 62.3 | 91.0 | 24.8∗ | 62.9 | 57.1 | 51.1∗ | 0.2 | |||
| temp | 41.9 | 87.2 | 26.0 | 56.3 | 58.5 | 57.4 | 5.4 | |||
Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | 74.8 | ||
| stop | 80.4 | 94.6 | 61.1 | 68.0 | 67.2 | 72.0 | 0.9 | |||
| temp | 64.5 | 84.1 | 54.0 | 58.0 | 56.8 | 54.3 | 12.8 | |||
Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | 70.4 | ||
| stop | 81.2 | 94.1 | 59.3 | 76.9 | 69.2 | 40.6 | 0.2 | |||
| temp | 74.1 | 80.0 | 46.5 | 30.6 | 58.3 | 25.4 | 17.9 | |||
Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | 81.6 | ||
| stop | 84.3 | 95.6 | 65.7 | 77.6 | 78.6 | 81.8 | 1.0 | |||
| temp | 76.6 | 93.9∗ | 61.2 | 72.7 | 70.3 | 59.6 | 9.2 | |||
Standard ICL | 70.7 | 85.0 | 93.2 | 58.3 | 66.7 | 66.3 | 55.0 | ||
| stop | 85.5 | 92.7 | 56.4 | 66.6 | 63.6 | 57.1 | 0.4 | |||
| temp | 69.8 | 82.8 | 56.3 | 58.8 | 67.5 | 42.9 | 7.7 | |||
Standard ICL | 73.6 | 82.8 | 94.9 | 62.8 | 74.6 | 71.2 | 55.4 | ||
| stop | 81.2 | 94.5 | 61.2 | 73.7 | 72.0 | 53.2 | 1.0 | |||
| temp | 71.1 | 95.6 | 61.0 | 72.4 | 72.9 | 54.3 | 2.4 | |||
Standard ICL | 80.2 | 82.2 | 97.0 | 67.4 | 82.4 | 73.6 | 78.6 | ||
| stop | 81.2 | 97.3 | 65.5 | 82.0 | 77.6 | 73.9 | 0.6 | |||
@@
Appendix K Comparison Experiments with More Text Examples
In the ideal scenario, our experiments would have been conducted on the full test set. However, in practice, this is infeasible for any of the models studied in our paper due to computational resource constraints. For instance, it took 42 hours for the OpenLlama 3B model to run one round of the representation ablation experiment on the whole test set of DBPedia (i.e., one cell in Table 3 for the DBPedia column). To verify our number of test examples decision, we provide additional results where we scale up the number of test examples and observe no difference with our original experimental setup. Thus, we believe that limiting our test set sample size to 500 is a reasonable setup. We provide the test set statistics and the experiment results in Table 14 and Table 15.
For TREC, RTE, and CB, using 500 test examples won’t affect the final results at all since their test set size is smaller than 500. We provide the results of experiments using all test examples in SST2, and 5000 test examples in AGNews and DBPedia here to prove our point that limiting our test set sample size to 500 is a reasonable compromise. Shown in Table 15, compared to the results we show in Table 3, the numbers are changed less than 1% for all the results.
Appendix L Results using different instruction prompts
@@
Appendix M Representation-level ablation on Machine Translation Tasks
Besides the classification tasks, we also show results in the machine translation tasks to show that our findings could also be extended in text generation tasks. We used the Flores MT dataset (costa2022no) to conduct this set of 4-shot machine translation experiments. The results are reported with the BLEU metric (papineni2002bleu). We investigated three different language directions: English-to-French, English-to-Danish, and English-to-German. We used 10 random seeds for En-Fr and En-De and 15 random seeds for En-Da to randomly choose the demonstrations. 100 test examples are sampled in this set of the experiments as a computational compromise. Similar to the classification tasks, we keep the answer (i.e., target language) unablated for all the settings and ablate different kinds of tokens. Results in Table 26 show the consistent finding to those in Section 5.2.
Appendix N Representation-level ablation on Question Answering Tasks
We show the representation-level experimental results of the question answering (QA) tasks in this section. We used Commonsense QA (talmor-etal-2019-commonsenseqa) dataset to test if the template and stopword tokens would directly affect the downstream task performance. We applied the settings of 4 in-context examples and 15 random seeds in this set of experiments. We frame the task as directly answering the questions instead of choosing one answer from the choices because the token types in this scenario are easier to be categorized.
@@
| Zero-shotcont | 7.42 | 16.62 | 14.49 | 19.47 | 17.51 | 17.78 | 16.64 | |||||||||||||||
| Zero-shotstop | 11.71 | 21.96 | 18.98 | 23.38 | 22.13 | 22.31 | 19.16 | |||||||||||||||
| Zero-shottemp | 13.24 | 24.38 | 25.73 | 27.20 | 25.42 | 25.11 | 25.56 | |||||||||||||||
| Standard ICLcont | 14.40 | 24.07 | 26.22 | 27.73 | 25.89 | 24.71 | 25.47 | |||||||||||||||
| Standard ICLstop | 11.96 | 21.84 | 21.44 | 26.93 | 24.62 | 23.09 | 23.69 | |||||||||||||||
@@
Appendix O Representation-level Ablation based on the Token Count
@@
| Tokenizer | Setting | Avg. | AGNews | DBPedia | SST2 | TREC | CB | RTE | ||
cont | 204.1 | 207.5 | 278.8 | 116.3 | 48.7 | 295.5 | 278.0 | ||
| stop | 43.0 | 43.2 | 45.1 | 27.7 | 21.8 | 66.7 | 53.7 | |||
| temp | 56.5 | 43.3 | 49.9 | 42.4 | 48.8 | 78.5 | 76.3 | |||
cont | 220.6 | 238.5 | 288.8 | 127.8 | 51.3 | 312.0 | 305.1 | ||
| stop | 45.0 | 43.1 | 46.1 | 29.9 | 21.7 | 71.1 | 57.9 | |||
| temp | 52.7 | 41.3 | 50.5 | 48.7 | 36.7 | 70.9 | 68.1 | |||
cont | 213.6 | 225.0 | 284.7 | 123.7 | 50.1 | 304.2 | 293.7 | ||
| stop | 44.7 | 43.2 | 45.0 | 29.9 | 21.7 | 70.7 | 57.8 | |||
@@
| Models | Setting | AGNews | DBPedia | SST2 | TREC | CB | RTE | Avg. | ||
| 10 Random Tokens | ||||||||||
From all | 14.5 | 16.9 | 71.7 | 47.3 | 60.2 | 57.5 | 44.7 | ||
| From cont | 3.0 | 9.2 | 60.9 | 61.9 | 65.2 | 60.4 | 43.4 | |||
| From stop | 15.2 | 35.6 | 75.2 | 68.5 | 62.0 | 64.1 | 53.4 | |||
| From temp | 53.0 | 50.1 | 59.1 | 56.2 | 64.4 | 60.2 | 57.2 | |||
From all | 67.8 | 67.2 | 75.0 | 70.9 | 63.8 | 60.7 | 67.6 | ||
| From cont | 36.6 | 64.0 | 63.1 | 67.5 | 59.3 | 57.3 | 58.0 | |||
| From stop | 73.5 | 68.6 | 86.3 | 72.8 | 61.8 | 61.6 | 70.8 | |||
| From temp | 78.7 | 78.7 | 92.1 | 68.9 | 71.2 | 68.4 | 76.3 | |||
| 20 Random Tokens | ||||||||||
From all | 12.6 | 24.0 | 77.2 | 54.4 | 60.6 | 58.9 | 47.9 | ||
| From cont | 1.3 | 8.6 | 65.3 | 63.5 | 61.7 | 62.2 | 43.8 | |||
| From stop | 28.3 | 45.5 | 84.5 | 66.8 | 58.0 | 67.1 | 58.4 | |||
| From temp | 75.0 | 63.7 | 60.3 | 59.0 | 65.4 | 60.7 | 64.0 | |||
From all | 78.4 | 67.8 | 74.0 | 70.1 | 65.4 | 62.0 | 69.6 | ||
| From cont | 69.1 | 63.3 | 59.3 | 68.2 | 60.5 | 56.8 | 62.9 | |||
| From stop | 78.4 | 74.3 | 89.9 | 74.2 | 68.3 | 69.1 | 75.7 | |||
| From temp | 81.9 | 78.7 | 93.0 | 68.9 | 72.4 | 71.4 | 77.7 | |||
| 30 Random Tokens | ||||||||||
From all | 27.4 | 27.7 | 77.9 | 54.9 | 63.9 | 62.2 | 52.3 | ||
| From cont | 1.3 | 7.6 | 77.9 | 66.3 | 64.2 | 63.5 | 46.8 | |||
| From stop | 44.8 | 59.8 | 92.8 | 68.3 | 53.2 | 69.3 | 64.7 | |||
| From temp | 76.9 | 66.1 | 61.0 | 59.6 | 67.9 | 59.8 | 65.2 | |||
From all | 79.9 | 74.6 | 83.0 | 67.8 | 69.0 | 64.6 | 73.2 | ||
| From cont | 71.0 | 62.6 | 58.2 | 67.2 | 59.2 | 59.4 | 62.9 | |||
| From stop | 79.8 | 76.9 | 91.0 | 74.2 | 71.1 | 68.8 | 77.0 | |||
| From temp | 84.0 | 79.3 | 93.8 | 68.9 | 75.6 | 75.4 | 79.5 | |||
| 40 Random Tokens | ||||||||||
From all | 45.2 | 27.2 | 77.9 | 54.9 | 58.9 | 62.3 | 54.4 | ||
| From cont | 0.9 | 9.1 | 89.5 | 66.1 | 62.3 | 62.4 | 48.4 | |||
| From stop | 49.2 | 63.7 | 94.0 | 68.7 | 53.7 | 70.8 | 66.7 | |||
| From temp | 77.9 | 66.1 | 62.9 | 61.5 | 68.1 | 60.1 | 66.1 | |||
From all | 78.8 | 73.4 | 86.9 | 69.2 | 69.3 | 67.4 | 74.2 | ||
| From cont | 74.1 | 59.4 | 55.8 | 68.0 | 61.0 | 59.9 | 63.0 | |||
| From stop | 80.1 | 76.4 | 91.0 | 74.2 | 72.0 | 69.6 | 77.2 | |||
| From temp | 85.0 | 80.7 | 93.8 | 69.0 | 76.4 | 76.2 | 80.2 | |||
Appendix P Results of the Token-level Ablation
@@
Detailed results of the token-level ablation are shown in Table 30. We omited the temp case from here since it constantly yields an accuracy of 0% when both the template token in the demonstrations and the test examples are ablated. Since the setting for the template tokens are not aligned with the ones for the stopword and content tokens, we included another set of experiments where only the the template tokens in the demonstrations are ablated at the token level in Appendix Q. We want to emphasize that this experimental design choice does not affect the main findings in Section 5.2, where information is being propagated from the representations of content tokens to the representations of the performance-critical tokens and this incorporation of the information makes them better at affecting task performance, partially explaining the working mechanism of in-context learning.
| Models | Settings | AGNews | SST2 | TREC | DBPedia | RTE | CB | Avg. | ||
Standard ICL | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | 58.0 | ||
| cont | 31.5 | 63.0 | 40.6 | 25.4 | 56.1 | 48.9 | 44.3 | |||
| stop | 64.4 | 91.5 | 20.9 | 62.3 | 57.8 | 52.6 | 58.3 | |||
Standard ICL | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | 74.8 | ||
| cont | 55.2 | 67.2 | 42.6 | 50.8 | 57.4 | 56.3 | 54.9 | |||
| stop | 82.3 | 93.8 | 64.1 | 69.7 | 66.5 | 70.0 | 74.4 | |||
Standard ICL | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | 70.4 | ||
| cont | 78.8 | 81.7 | 45.3 | 75.1 | 55.1 | 54.5 | 65.1 | |||
| stop | 82.5 | 92.5 | 61.5 | 76.5 | 69.6 | 40.5 | 70.5 | |||
Standard ICL | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | 81.6 | ||
| cont | 74.0 | 89.6 | 67.0 | 73.0 | 69.8 | 49.0 | 70.4 | |||
@@
Appendix Q Token-level Ablation for Template Tokens
To maintain consistency of the templates across both demonstrations and test examples, we choose to ablate the template tokens at the token level in both in Section 5.2. This experimental design differs from the other two token-level ablations. This inconsistency does not impact the main findings in Section 5.2, which show that information is propagated from the representations of content tokens to the representations of performance-critical tokens and this information aggregation enhances the ability of performance-critical tokens to improve the final task performance, partially explaining the mechanism of in-context learning. For completeness, we provide a supplemental experiment in this section where only the template tokens in the demonstrations are ablated.
@@
| Models | Settings | Avg. | AGNews | SST2 | TREC | DBPedia | RTE | CB | ||
Standard ICL | 58.0 | 63.7 | 91.2 | 21.9 | 61.9 | 57.4 | 52.0 | ||
| temp | 17.4 | 0.0 | 24.2 | 40.5 | 0.4 | 35.6 | 3.6 | |||
Standard ICL | 74.8 | 82.4 | 94.3 | 63.5 | 68.7 | 68.6 | 71.3 | ||
| temp | 29.0 | 41.0 | 11.4 | 39.9 | 0.9 | 49.7 | 31.3 | |||
Standard ICL | 70.4 | 81.6 | 94.3 | 60.0 | 76.1 | 70.6 | 39.9 | ||
| temp | 36.9 | 82.1 | 17.5 | 51.4 | 8.6 | 39.5 | 22.0 | |||
Standard ICL | 81.6 | 85.0 | 96.5 | 68.1 | 78.4 | 78.5 | 83.3 | ||
@@
Appendix R Possible Applications
@@
This area is indeed attracting increased research attention, and our findings could contribute valuable insights into ongoing work on efficient sequence modeling and memory management in LLMs (NEURIPS2023_a452a7c6; zhang2023h_2o; bai2024citrus). By identifying which tokens retain critical task-related information, our work aligns with and can inform methods focused on compressing intermediate states and improving long-context processing for applications ranging from summarization and document understanding to interactive dialogue systems.
@@
Given our findings, there is a promising avenue for improving in-context learning (ICL) performance by including a greater number of examples in ICL prompts (li2023incontext; hao2022structured; bertsch2024incontext). Our results suggest that only a subset of token representations, specifically performance-critical tokens, play a critical role in determining ICL performance, while the representations of other tokens are less impactful. This observation opens up the possibility of selectively compressing or omitting unimportant token representations after the initial encoding of a demonstration. By doing so, it becomes feasible to maximize the use of the model’s fixed-length capacity, potentially enabling the inclusion of a higher number of examples within the same prompt length constraints. This approach may enhance the effectiveness of ICL in tasks where the availability of diverse examples contributes to improved model accuracy and stability.
@@
Appendix S Results of Representation-level Partial performance-critical Token Ablation